A solenoid of radius 2.60 cm has 350 turns and a length of 15.0 cm.

(a) Find its inductance.
=___mH

(b) Find the rate at which current must change through it to produce an emf of 72.0
Answer #1

a) Inductance of solenoid, L = mue*N^2*A/l

= 4*pi*10^-7*350^2*(pi*0.026^2)/0.15

= 2.18*10^-3 H

= 2.18 mH

b) Induced emf = L*dI/dt

==> dI/dt = induced emf/L

= 72*10^-6/(2.18*10^-3)

= 33 mA/s